OpenStudy (anonymous):

Find the discriminant of the function f(x) = x2 - 4x + 5 and determine how many roots the function has.

OpenStudy (tommo_lcfc):

x^2 - 4x + 5 The discriminant of this function is b^2 - 4ac which in this case is (-4)^2 - 4(1)(5) = 16 - 20 = -4. Therefore, the function does not have any REAL roots, but it does have complex roots. The complex roots of this function are -2 + 2i and -2 - 2i (using the quadratic formula). Hope this helps :)